# Redistricting Reforms Reduce Gerrymandering by Constraining Partisan Actors

### v1

@@ 4 Estimating the Causal Effects of Institutional Leeway
Thus, we assume that once we adjust for the state-specific change in political geography, states with different changes in institutional features have parallel trends in the potential outcomes. Formally, if we denote the difference between the potential outcome and the simulated outcome as , our conditional parallel trend assumption becomes
@@
We estimate this causal estimand with a Bayesian linear regression model, where the response is the change in each simulation-adjusted outcome between 2010 and 2020, . Given the small sample size () and the moderate number of covariates ( with interactions), a Bayesian model is useful, as the priors help avoid overfitting to a few samples, and uncertainty is automatically quantified. A more flexible regression model beyond linear would be unlikely to increase predictive power, given the small sample size, and would further risk overfitting. The predictors are the change in treatment level , the baseline treatment level , the covariates , and the interaction of treatment change with baseline treatment and with each of the covariates.
@@ 5.1 Outcome measures
Our measures can be divided into two buckets: nonpartisan outcomes that quantify how much gerrymandering is present or how responsive a districting plan is to shifts in public opinion, and partisan outcomes that capture which party is advantaged by a districting plan. For the nonpartisan outcomes, we also use the absolute value of the realized leeway as an additional treatment variable, so that neither treatment nor outcome consider the partisan direction.
@@
We adopt the model from Kenny et al., (2023), which assumes district election outcomes can be decomposed into a baseline district-level vote share plus district-specific and national swings.121212The full details of this approach can be found in the Methods section and SI Appendix, Section B of Kenny et al., (2023). The model is closely related to the stochastic uniform partisan swing model of Gelman and King, (1994) and the congressional model of Ebanks et al., (2023). As the baseline district-level vote, we use the 2008 presidential results for the 2010 redistricting cycle and an average of the 2016 and 2020 presidential results for the 2020 cycle, each logit-shifted so that the national partisan vote is exactly 50/50. As in Kenny et al., (2023), we use historical congressional election data since 1976 to estimate the variance of the district and national swings. We calculate exact expectations of all of our outcome measures against the predictive distribution of the model via numerical integration.
Our primary partisan measure of districting plans is the expected number of seats won by the Republican party. As discussed in Section 4, we subtract the average outcomes in the simulated baseline redistricting plans (McCartan et al.,, 2022; Kenny et al.,, 2024) from the observed outcome in the enacted plan to adjust for the state-specific change in political geography, which is a potential confounding factor. After this adjustment, the seat outcome ranges from –1.57 in Illinois in 2020 (favoring Democrats) to 1.78 in Texas in 2020 (favoring Republicans). These simulation-differenced seats can be directly interpreted as a measurement of bias due to partisan gerrymandering, with positive values indicating a Republican bias beyond what would be expected based on the state’s political geography alone.
Republican seats is an interpretable measure, but may not be completely comparable across states. Depending on political geography and especially the total number of districts in each state, the natural variation in the number of seats outcome may vary significantly between states. To address this, we also include as an outcome measure the simulation -score of Republican seats, which is calculated by taking the simulation-differenced seats outcome and dividing it by the standard deviation of Republican seats in the simulation set (Kenny et al.,, 2023). This puts all the states’ outcomes on a common scale, increasing the plausibility of the parallel trends assumption and the homoskedasticity assumption of the estimation model.
@@
We also study two more outcome measures based on the efficiency gap, a popular measure of partisan bias that is calculated as the difference between the two parties’ wasted votes (Stephanopoulos and McGhee,, 2015). Wasted votes are votes cast for losing candidates and votes cast for winning candidates beyond the threshold needed to win. The idea behind the efficiency gap is that a fair plan should have similar numbers of wasted votes for both parties. The partisan-signed version of the efficiency gap indicates which party benefits from the plan in terms of vote efficiency; we use positive values to indicate pro-Republican bias. Unlike with Republican seat share, we do not subtract the average simulated value from the partisan-signed efficiency gap measure. This is because the efficiency gap is already designed to have a universal scale, where 0 indicates partisan fairness; subtracting the simulated average would change the definition of a fair baseline for this measure. To create a nonpartisan measure, as with Republican seats we take the absolute value of the difference between the enacted and average simulated plan’s efficiency gap. Smaller values indicate a fairer plan by this measure (closer to the normatively preferred value of 0).
Finally, we also measure the responsiveness of districting plans to changes in the national electoral environment. Responsiveness is measured as the rate of change in the share of Republican seats given an infinitesimal change in Republican vote share nationwide. Responsiveness is closely linked to the presence of competitive seats: the more competitive seats there are, the larger the change in seat share will be for a given shift in vote share. Indeed, electoral responsiveness is the primary motivation for the creation of competitive congressional districts.
The responsiveness measure ranges from 0.044 in Wyoming in 2020 to 7.91 in New Hampshire in 2020, with most plans’ values lying between 0.5 and 3. The interpretation of these values is as follows: in New Hampshire, a 1pp increase in a party’s vote share leads to a 7.91pp change in the party’s expected seat share. This relatively large increase makes sense in context—both of the state’s congressional districts have Republican vote share within a few points of 50%. As with the other measures, we subtract the mean responsiveness of the simulated plans from the enacted plan’s responsiveness when estimating causal effects.
@@
5.2 Results
We first study the effects of changes in maximum leeway on Republican seats. Recall that maximum leeway represents the worst expected bias under a single party control. The left panel of Figure 6 shows the estimated coefficients of the estimation model discussed in Section 4 fit to the simulation-adjusted seats variable. All coefficients are shown on the scale of the outcome, where a positive coefficient indicates a relationship with a positive change in Republican seats from 2010 to 2020. The model’s is around 0.4, highlighting the potential for the control variables to confound any observed correlation between reform and the outcome measure.
Given the fitted outcome model, we can estimate the conditional average treatment effect (CATE) for any given starting and final value of maximum leeway and any combination of covariates. The right panel of Figure 6 plots these CATEs for each state that experienced a change in maximum leeway between 2010 and 2020, using that state’s specific covariates. There is a clear dose-response relationship between the decrease in maximum leeway and the decrease in the expected number of Republican seats.141414This pattern was also observed in robustness checks that used a flexible Bayesian Additive Regression Trees (BART) model for estimation. A state with Michigan’s dose (nearly 4) has an estimated CATE of around –1.25, with 95% credible interval (–2.1, –0.6), and so would be expected to gain just over 1 Democratic seat, on average. Similarly, a state with Ohio’s dose (just over 1) and covariates would be expected to gain about half of a Democratic seat (0.2, 0.6). Motivated by this pattern, we also include in the right figure a dose-response curve calculated by calculating a CATE for dose level and averaging over the observed covariate distribution.151515Because of the interaction between the dose variable and the covariates (which do not have mean zero), the overall slope estimate and its variance differ from the estimate and variance of the coefficient on Dose in the linear model.
Since the dose-response pattern is well-summarized by a linear relationship, we can summarize it by its slope, a quantity also known as the average causal response (ACR). The product of the ACR and a given dose correspond to the estimated causal effect of that dose, averaged across states.
@@ 5.2 Results
Across the three nonpartisan outcome measures measuring the amount of gerrymandering, we find a consistent effect of leeway. The estimated ACRs are all positive, meaning an increase in leeway leads to an increased gap between the enacted plan and the simulated baseline; equivalently, states that reduce leeway through reform efforts are expected to rein in partisan gerrymandering. The magnitude of the effect is substantial: a reform effort like Michigan’s, with a decrease in maximum leeway of 4 units, would be expected to reduce the difference in Republican seats from the baseline by about 0.8 seats (0.2, 1.3), or 3.3 simulation standard deviations (–0.1, 6.8)—enough to move an extreme partisan gerrymander to a fair plan. The ACRs for absolute realized leeway are very similar to those for maximum leeway; the ACRs for the fully signed realized leway are positive but smaller in magnitude, and have enough uncertainty that the 80% credible intervals cross zero. Thus there is some weak evidence that giving Democrats more control of the redistricting process by increasing their leeway or reducing Republican leeway would also reduce gerrymandering.
@@
We also find that reforms increase electoral responsiveness. The point estimates of the ACRs for both maximum leeway and realized leeway agree; the posterior probability that the effect is negative is 88% and 95%, respectively. There estimated effect for absolute realized leeway is near zero. We estimate that a reduction in maximum or realized leeway of 4 units would increase responsiveness by 0.61 (–0.39, 1.66), meaning that a 1pp increase in a party’s vote share would translate to an additional 0.6pp in seat share on top of the average 2pp increase in seat share across all states.
@@
Another way to interpret the responsiveness effect is in terms of the share of competitive seats, where a seat that is counted as competitive in proportion to how close its baseline partisanship is to 50%. It turns out that this is equivalent to a linear rescaling of the responsiveness outcome.161616Specifically, we measure the competitiveness of a single seat as the derivative of the election model’s win probability for that seat at its baseline vote, divided by the maximum derivative, which is obtained at a 50/50 baseline vote. Based on our election model, a district with 50/50 baseline vote is counted as 1.0 competitive seats, and a district with 60/40 baseline vote is counted as just 0.14 competitive seats. This approach avoids arbitrary cutoffs in what counts as a competitive district. In terms of competitive seats, we estimate that a reduction in leeway of 4 units would increase the share of competitive seats in an average state from 25% to 32% (20%, 45%).
@@
In terms of partisan outcomes, the estimated ACRs on Republican seats are all positive, for both realized and maximum leeway. This is in line with an interpretation that increasing leeway benefits the party in control, as one might expect. Recall that the doses from Figure 4, for the realized leeway estimates, a 4 dose is approximately going from a Democratic legislature to an independent commission or an independent commission to a Republican legislature. For such a dose of 4, we would expect Republicans to gain around 0.69 seats (0.21, 1.15), or 7 simulation standard deviations (2.4, 11.2)—again, a substantial effect. We do not find a consistent effect on the signed efficiency gap outcome, which was not adjusted using the simulations.
@@ 6 Redistricting Commissions: Counterfactual Policy Analysis
In this section, we conduct a counterfactual policy analysis about the nationwide adoption of redistricting commissions. Specifically, we use our causal models to predict the electoral outcomes under a hypothetical scenario where every state adopts a particular type of redistricting commissions. Although each of the 15 states uses a redistricting commission with different structures and rules, our model is able to characterize these different commissions in terms of their partisan leeway, and use these treatment values to predict electoral outcomes under a given commission structure.
@@
We study three types of commission structures currently enacted in several states: (1) a New York-style commission with a nonpartisan map drawer and several partisan veto points; (2) an Ohio-style reform with legislature-drawn map and several partisan and bipartisan veto points and stalemate procedures, including a bipartisan commission stalemate procedure; (3) and a Michigan-style reform, with a nonpartisan commission, no partisan veto points, and the potential for court review.
Figure 8 shows a seats-votes curve for each counterfactual commission scenario (the orange lines), which display the predicted number of total Democratic seats in the US House (y-axis) for a given national Democratic vote share (x-axis).
Seats-votes curves are often used to study a measure of partisan bias, operationalized as the difference in the share of seats and votes at a given point on the curve (Tufte,, 1973; King and Browning,, 1987; Katz et al.,, 2020). A seats-votes curve for which a 50% vote share translates exactly to a 50% seat share represents the baseline for partisan bais. The lines for 2010 and 2020 intersect the 50% vote line below the 50% seat line, since a combination of the underlying geographic distribution of Democratic votes and gerrymandering efforts nationwide disadvantages the Democratic party.
@@
We approximate the seats-votes curve around 50% vote share as a line, and estimate the effect of reforms on its slope and intercept separately. Up to a linear transformation, the slope corresponds to responsiveness and the intercept to seat share. These two outcome variables are differenced using the simulation sets, as discussed in Section 4.
We find that nationwide adoption of commissions would generally reduce partisan bias, but the extent of this reduction largely depends on details of commission structure. The model predicts that all commission structures would reduce the existing Republican advantage in the House driven by both the inefficient geographic distribution of Democratic voters (Chen and Cottrell,, 2016; Chen and Rodden,, 2015) and net partisan manipulation favoring Republicans (Kenny et al.,, 2023).
@@
We find that Michigan-style reforms have the greatest effects in both responsiveness and Democratic seats, since partisan actors are the most constrained by the presence of a nonpartisan commission, no partisan veto points, and the potential for court review. We estimate that the number of Democratic seats gained under Michigan-style reforms is 22 with 95% credible interval (7.4, 38.3), and that for every 1pp increase in vote share, Democrats gain an estimated 9 seats (6.8, 11.2), an increase of 1.2 over the baseline. We are more than 99% confident that the Michigan-style reforms have a larger effect than the Ohio reforms, and more than 99% confident in a larger effect than the New York-style reforms. We are also 87% confident that the Michigan-style reform increases responsiveness.
Though all three proposed reforms reduce the deviation from partisan symmetry by increasing net Democratic seats, Michigan-style reforms suggest that much greater effects are produced when multiple nodes of the redistricting game constrain leeway by ensuring partisan nodes are nonpartisan, and do not precede partisan vetoes. Meanwhile, the examples of New York-style and Ohio-style reforms demonstrate how constraining partisan actors at different nodes, through different reforms, may produce substantively similar effects.

### v2

@@ 4.2 Adjusting for political geography
One limitation of basing simulations off of each state’s rules is that any changes in outcome variables from 2010 to 2020 that are due to changes in rules would be subtracted away. Since redistricting rules are sometimes changed as part of larger reform efforts, part of the overall effect of redistricting reform would not be included in the measured effect, attenuating the overall estimated effects. However, these rules changes occurred in only a few states, and their effect on partisan outcomes was measured by Kenny et al., (2023) and found to be minimal.
Thus, we assume that once we adjust for the state-specific change in political geography, states with different changes in institutional features have parallel trends in the potential outcomes. Formally, if we denote the difference between the potential outcome and the simulated outcome as , our conditional parallel trends assumption becomes
@@
4.3 Estimation
[…]
@@ 5.1 Outcome measures
Our measures can be divided into two buckets: nonpartisan outcomes that quantify how much gerrymandering is present or how responsive a districting plan is to shifts in public opinion, and partisan outcomes that capture which party is advantaged by a districting plan. For the nonpartisan outcomes, we also use the absolute value of the realized leeway as an additional treatment variable so that neither treatment nor outcome considers the partisan direction.
@@
We adopt the model from Kenny et al., (2023), which assumes district election outcomes can be decomposed into a baseline district-level vote share plus district-specific and national swings.121212The full details of this approach can be found in the Methods section and SI Appendix, Section B of Kenny et al., (2023) The model is closely related to the stochastic uniform partisan swing model of Gelman and King, (1994) and the congressional model of Ebanks et al., (2023). As the baseline district-level vote, we use the 2008 presidential results for the 2010 redistricting cycle and an average of the 2016 and 2020 presidential results for the 2020 cycle, each logit-shifted so that the national partisan vote is exactly 50/50 (Voting and Election Science Team,, 2018, 2020). As in Kenny et al., (2023), we use historical congressional election data since 1976 to estimate the variance of district and national swings. We calculate exact expectations of all of our outcome measures against the predictive distribution of the model via numerical integration.
Our primary partisan measure of districting plans is the expected number of seats won by the Republican party. As discussed in Section 4, we subtract the average outcomes in the simulated baseline redistricting plans (McCartan et al.,, 2022; Kenny et al.,, 2024) from the observed outcome in the enacted plan to adjust for the state-specific change in political geography, which is a potential confounding factor (Cottrell,, 2019). After this adjustment, the seat outcome ranges from in Illinois in 2020 (favoring Democrats) to in Texas in 2020 (favoring Republicans). These simulation-differenced seats can be directly interpreted as a measurement of bias due to partisan gerrymandering, with positive values indicating a Republican bias beyond what would be expected based on the state’s political geography alone.
The expected number of Republican seats is an interpretable measure, but may not be completely comparable across states. Depending on political geography and especially the total number of districts in each state, the natural variation in the number of seats outcome may vary significantly between states. To address this, we also include as an outcome measure the simulation -score of Republican seats, which is calculated by taking the simulation-differenced seats outcome and dividing it by the standard deviation of Republican seats in the simulation set (Kenny et al.,, 2023). This puts all the states’ outcomes on a common scale, increasing the plausibility of the parallel trends assumption and the homoskedasticity assumption of the estimation model.
@@
We also study two more outcome measures based on the efficiency gap, which is a popular measure of partisan bias calculated as the difference between the wasted votes of the two parties (Stephanopoulos and McGhee,, 2015). Wasted votes are votes cast for losing candidates and votes cast for winning candidates beyond the threshold needed to win. The idea behind the efficiency gap is that a fair plan should have similar numbers of wasted votes for both parties. The partisan-signed version of the efficiency gap indicates which party benefits from the plan in terms of vote efficiency; we use positive values to indicate pro-Republican bias. Unlike with Republican seat share, we do not subtract the average simulated value from the partisan-signed efficiency gap measure. This is because the efficiency gap is already designed to have a universal scale, where 0 indicates partisan fairness; subtracting the simulated average would change the definition of a fair baseline for this measure. To create a nonpartisan measure, as with Republican seats, we take the absolute value of the difference between the enacted and average simulated plan’s efficiency gap. Smaller values indicate a fairer plan by this measure (closer to the normatively preferred value of 0).
Finally, we also measure the responsiveness of districting plans to changes in the national electoral environment. Responsiveness is measured as the rate of change in the share of Republican seats given an infinitesimal change in Republican vote share nationwide. Responsiveness is closely linked to the presence of competitive seats; the more competitive seats there are, the larger the change in seat share will be for a given shift in vote share. Indeed, electoral responsiveness is the primary motivation for the creation of competitive congressional districts.
The responsiveness measure ranges from 0.044 in Wyoming in 2020 to 7.91 in New Hampshire in 2020, with most plans’ values lying between 0.5 and 3. The interpretation of these values is as follows; in New Hampshire, a 1pp increase in a party’s vote share leads to a 7.91pp change in the party’s expected seat share. This relatively large increase makes sense in context—both of the state’s congressional districts have Republican vote share within a few points of 50%. As with the other measures, we subtract the mean responsiveness of the simulated plans from the enacted plan’s responsiveness when estimating causal effects.
@@
5.2 Findings
[…]
Given the fitted outcome model, we can estimate the conditional average treatment effect (CATE) for any given starting and final value of maximum leeway and any combination of covariates. The right panel of Figure 5 plots these CATEs for each state that experienced a change in maximum leeway between 2010 and 2020, using the specific covariates of that state. There is a clear dose-response relationship between the decrease in maximum leeway and the decrease in the expected number of Republican seats.141414This pattern was also observed in robustness checks that used a flexible Bayesian Additive Regression Trees (BART) model for estimation. A state with Michigan’s dose (nearly 4) has an estimated CATE of around , with 95% credible interval (–2.0, –0.4), and so would be expected to gain just over 1 Democratic seat, on average. Similarly, a state with Ohio’s dose (just over 1) and covariates would be expected to gain about half of a Democratic seat (–0.6, –0.1). Motivated by this pattern, we also include in the right figure a dose-response curve calculated by calculating a CATE for dose level and averaging over the observed covariate distribution.151515Because of the interaction between the dose variable and the covariates (which do not have mean zero), the overall slope estimate and its variance differ from the estimate and variance of the coefficient on Dose in the linear model.
Since the dose-response pattern is well approximated by a linear relationship, we can summarize it by its slope, a quantity also known as the average causal response (ACR). The product of the ACR and a given dose corresponds to the estimated causal effect of that dose, averaged across states.
@@ 5.2 Findings
Across the three nonpartisan outcomes measuring the amount of gerrymandering, we find a consistent effect of leeway. The estimated ACRs are all positive, meaning that an increase in leeway leads to a greater gap between the enacted plan and the simulated baseline. That is, states that reduce leeway through reform efforts are expected to rein in partisan gerrymandering. The magnitude of the effect is substantial: a reform effort like Michigan’s, with a decrease in the maximum leeway of 4 units, would be expected to reduce the difference in Republican seats from the baseline by about 0.5 seats (–0.006, 1.040), or 1.7 simulation standard deviations (–2, 5)—enough to move an extreme partisan gerrymander to a fair plan. The ACRs for the absolute realized leeway are very similar to those for the maximum leeway; the ACRs for the fully signed realized leeway are positive but smaller in magnitude, and have enough uncertainty that the 80% credible intervals cross zero. Thus, there is some weak evidence that giving Democrats more control of the redistricting process by increasing their leeway or reducing Republican leeway would also reduce gerrymandering.
@@
We also find that reforms increase electoral responsiveness. The point estimates of the ACRs for both the maximum leeway and the realized leeway agree; the posterior probability that the effect is negative is 97% and 96%, respectively. The estimated effect for the absolute realized leeway is near zero. We estimate that a reduction in the maximum or realized leeway of 4 units would increase responsiveness by 1.1 (–0.042, 2.142), meaning that a 1pp increase in a party’s vote share would translate to an additional 0.6pp in seat share on top of the average 2pp increase in seat share across all states.
@@
Another way to interpret the effect on responsiveness is in terms of the share of competitive seats, where a seat that is counted as competitive in proportion to how close its baseline partisanship is to 50%. This is equivalent to a linear rescaling of the responsiveness outcome.161616Specifically, we measure the competitiveness of a single seat as the derivative of the election model’s win probability for that seat at its baseline vote, divided by the maximum derivative, which is obtained at a 50/50 baseline vote. Based on our election model, a district with 50/50 baseline vote is counted as 1.0 competitive seats, and a district with 60/40 baseline vote is counted as just 0.14 competitive seats. This approach avoids arbitrary cutoffs in what counts as a competitive district. In terms of competitive seats, we estimate that a reduction in leeway of 4 units would increase the share of competitive seats in an average state from 25% to 38% (24%, 51%).
@@
In terms of partisan outcomes, the estimated ACRs on Republican seats are all positive, for both realized and maximum leeway. This is in line with an interpretation that increasing leeway benefits the party in control, as one might expect. Recall that the doses from Figure 3, for the realized leeway estimates, a dose of 4 represents approximately going from a Democratic legislature to an independent commission or an independent commission to a Republican legislature. For such a dose of 4, we would expect Republicans to gain around 0.7 seats (0.21, 1.18), or 7.2 simulation standard deviations (2.4, 12.2)—again, a substantial effect. We do not find a consistent effect on the signed efficiency gap outcome, which was not adjusted using the simulations.
@@ 6 Redistricting Commissions: Counterfactual Policy Analysis
In this section, we conduct a counterfactual policy analysis of the nationwide adoption of redistricting commissions. Though hypothetical, counterfactual policy analyses are useful tools for evaluating institutional reform proposals (Cervas and Grofman,, 2019). Our approach is well suited for this analysis, in part because commissions can vary drastically in the extent to which they limit the influence of partisan actors. We use our causal models to predict electoral outcomes under a series of hypothetical scenarios, in which every state adopts a redistricting commission with particular institutional powers. Although each of the 15 states uses a redistricting commission with different structures and rules, our model is able to characterize these different commissions in terms of their partisan leeway and use these treatment values to predict electoral outcomes under a given commission structure.
@@
We study three types of commission structures currently enacted in several states: (1) a New York-style commission with a nonpartisan map drawer and several partisan veto points; (2) an Ohio-style reform with legislature-drawn map and several partisan and bipartisan veto points and stalemate procedures, including a bipartisan commission stalemate procedure; (3) and a Michigan-style reform, with a nonpartisan commission, no partisan veto points, and the potential for court review.181818Although states could change their redistricting processes in the future, we use shorthands like “Michigan-style” to refer to the current (as of 2024) institutional design of each state’s commission structure. Though each of these states has a commission of some kind, they differ in important ways — for example, Michigan’s elimination of partisan veto points removes the ability of partisan actors to veto unfavorable plans.
@@
We visualize the impact of each reform structure using a seats-votes curve. Seats-votes curves are often used to study a measure of partisan bias, operationalized as the difference in the share of seats and votes at a given point on the curve (Tufte,, 1973; King and Browning,, 1987; Katz et al.,, 2020). A seats-votes curve for which a 50% vote share translates exactly to a 50% seat share represents the baseline for partisan bias. We approximate the seats-votes curve around 50% vote share as a line, and estimate the effect of reforms on its slope and intercept separately. Up to a linear transformation, the slope corresponds to responsiveness and the intercept to seat share. These two outcome variables are differenced using the simulation sets, as discussed in Section 4.
We find that nationwide adoption of commissions would generally reduce partisan bias, and that reforms that place additional restrictions on partisan actors (as in Michigan) are generally more effective. Figure 7 shows a seats-votes curve for each counterfactual commission scenario (the orange lines), which display the predicted number of total Democratic seats in the US House (y-axis) for a given national Democratic vote share (x-axis). The model predicts that all commission structures would reduce the existing Republican advantage in the House driven by both the inefficient geographic distribution of Democratic voters (Chen and Cottrell,, 2016; Chen and Rodden,, 2015) and net partisan manipulation favoring Republicans (Kenny et al.,, 2023). We also compare these estimated relationships to the actual seats-votes curves for 2010 and 2020 (the black lines). We note that the lines for 2010 and 2020 intersect the 50% vote line below the 50% seat line, since a combination of the underlying geographic distribution of Democratic votes and gerrymandering efforts nationwide disadvantages the Democratic party.
@@
Further, we find that Michigan-style reforms have the greatest effects in both responsiveness and Democratic seats, since partisan actors are the most constrained by the presence of a nonpartisan commission, no partisan veto points, and the potential for court review. We estimate that the number of Democratic seats gained under Michigan-style reforms is 22 with 95% credible interval (7.4, 38.3), and that for every 1pp increase in vote share, Democrats gain an estimated 9 seats (6.8, 11.2), an increase of 1.2 over the baseline. We are more than 99% confident that Michigan-style reforms have a larger effect than Ohio reforms, and more than 99% confident in a larger effect than New York-style reforms. We are also 87% confident that Michigan-style reforms increase responsiveness.
Though all three proposed reforms reduce the deviation from partisan symmetry by increasing net Democratic seats, Michigan-style reforms suggest that much greater effects are produced when multiple nodes of the redistricting game constrain leeway by ensuring partisan nodes are nonpartisan, and do not precede partisan vetoes. Meanwhile, examples of New York-style and Ohio-style reforms demonstrate how constraining partisan actors at different nodes, through different reforms, may produce substantively similar effects.